Circular No. 7226

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1999 N4 (LINEAR) Another LINEAR discovery with unusual motion placed in The NEO

Confirmation Page has been found to have a retrograde orbit:

1999 UT  α_{2000}  δ_{2000}  m_{2}
July 12.25443  19^{h}13^{m}44.^{s}43  +1^{o}07′42′′.1  18.2

At the request of the Central Bureau, some of the observers making astrometric observations examined their images carefully and have concluded that the object is a comet. M. Tichý (Kleť, 0.6-m reflector) noted comae of diameter ∼ 6′′ on July 14.9 and ∼ 7′′ on July 15.9 UT. L. Šarounová (Ondřejov, 0.6-m reflector) indicated a small coma, some lack of condensation but no tail on July 16.9. F. B. Zoltowski (Woomera, 0.3-m reflector) remarked on a compact, diffuse image with an asymmetrical distribution that may indicate a small, faint tail at p.a. ∼ 90^{o} on July 17.6.

Further astrometric observations, the following preliminary (and rather rough) parabolic orbital elements and an ephemeris are given on MPEC 1999-O06:

T = 2000 May 3.239 TT  ω = 87.^{o}861
Ω = 345.805 }2000.0
q = 5.57961 AU  i = 156.895

V382 VELORUM A. C. Gilmore, Mount John University Observatory, provides further

photometric observations, continuing the series on IAUC 7216: July 8.449 UT, V = 7.50, U − B = −0.89, B − V = −0.17, V − R = +1.33, V − I = +0.74, airmass = 2.00; 9.386, 7.50, −0.91, −0.19, +1.29, +0.68, 1.51; 13.418, 7.53, −0.90, −0.18, +1.25, +0.65, 1.83; 14.336, 7.61, −0.91, −0.20, +1.36, +0.67, 1.32. Standard deviations are generally < 0.01 mag.

COMET C/1995 O1 (HALE-BOPP) Total visual magnitude estimates: Mar. 17.46 UT, 11.2 (D. A. J. Sear-

gent, Cowra, N.S.W., 25 × 100 binoculars); Apr. 4.47, 11.5 (A. Pearce, Guilderton, Western Australia, 0.20-m reflector); 14.47, 11.8 (S. T. Rae, Palmerston North, New Zealand, 0.25-m reflector); 21.59, 11.9 (M. Mattiazzo, Wallaroo, South Australia, 0.20-m reflector); May 8.37, 12.2 (Rae); 22.32, 12.0 (Rae); June 2.43, 12.4 (Mattiazzo); 7.30, 12.1 (Rae); 15.40, 11.5 (Seargent); 19.75, 12.4 (Rae); July 9.41, 12.4 (Mattiazzo).

1999 July 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Brian G. Marsden